OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (anonymous):

$81.09

OpenStudy (paxpolaris):

\[C = 0.765x + 0.06\left( 0.765x \right) \]\[=1.06 \times \left( 0.765\times x \right)\]\[=1.06\times 0.765 \times 100\]\[=106\times0.765\] \[\huge \begin{matrix}\ \ \ \ 0.765 \\ \times\ \ \ 106 \\ \overline{\ \ \ 4590}\\ 765\ \ \ \\ \overline{\ 81.090}\end{matrix}\]